Skip to end of metadata
Go to start of metadata

Still trying to work several problems:

  1. P.0m is only 1Hz, only about same resolution as Pref, so not much help monitoring Pref after all.  Thus, change back to monitor Pref with 11b's 202BG
  2. Data from P.0m yesterday showed Pref fluctuations.  Try to damp these by removing potential leak (old capillary leak) and add more steel wool.

~0900: opened ref volume

  • - removed cap with capillary leak
  • - stuffed more steel wool (about double) in this side
  • - replace cap with solid cap
  • - removed sample cap
  • - stuffed more steel wool in this side
  • - replaced cap

0925-0937: sat in this configuration to allow P.0m to monitor Pref

0937:

  • - removed garden hose from CHATS system, noted that big hose fitting was somewhat loose, repaired with teflon tape?
  • - changed P.0m from AHATS Pref to garden hose/Bedard
  • - plumbed Pref back to CHATS system (using 11b.4.8m 202BG)

0955: covered up with dirt/space blanket again